OpenStudy (anonymous):

4x^(-2/3) Help me simplify please? :)

OpenStudy (anonymous):

So far, I've gotten to (1/4x)^2/3

OpenStudy (anonymous):

Then, I thought maybe the simplification would go to the following in the photo, but i'm not sure if that is it

OpenStudy (xapproachesinfinity):

recall that \(\large A^{-n}= \frac{1}{\large A^n}\)

OpenStudy (xapproachesinfinity):

and recall that \(\large \rm \color{green}{\sqrt[n]{x^m}=x^{\frac{m}{n}}}\)

OpenStudy (xapproachesinfinity):

that's pretty much what you need to do this